Bailand Company purchased a building for $286,000 that had an estimated residual value of $6,000 and an estimated service life of 10 years. Bailand purchased the building 4 years ago and has used straight-line depreciation. At the beginning of the fifth year (before it records depreciation expense for the year), the following independent situations occur: 1. Bailand estimates that the asset has 8 years’ life remaining (for a total of 12 years). 2. Bailand changes to the sum-of-the-years’-digits method. 3. Bailand discovers that the estimated residual value has been ignored in the computation of depreciation expense. Required: For each of the independent situations, prepare all the journal entries relating to the building for the fifth year. Ignore income taxes.
Answer #1

Given:

Bailand Company purchased a building for $286,000

Building's estimated residual value of $6,000

It's estimated service life of 10 years.

Bailand purchased the building 4 years ago and has used straight-line depreciation.

Depreciated cost = (286,000-6,000)/10 = $28,000

Amount Depreciated in the last 4 years = 4*28,000 = $ 112,000

1) If the life time of the building is 12 years, Since we have paid the depreciation of $112,000 for the last 4 years we will calculate the remaining amount for the next 8 years through straight line method using life time = 8 years

Depreciation cost = (280,000-112,000)/8 = $21,000

Journal Entry Credit Debit
Accumulated Depreciation expense Account $28,000
Long term Asset account $28,000
Accumulated Depreciation expense account $7000
Adjusted Depreciation expense $7000
Adjusted Depreciation expense account $7000
Long term Asset account $7000

2) The remaining depreciation amount = 280000-112000 = $168,000

For the next 6 years the sum of years = 1+2+3+4+5+6 = 21

depreciation cost = 168000* 6/21 = $48,000

Journal Entry Credit Debit
Accumulated Depreciation expense Account $28,000
Long term Asset account $28,000
Adjusted Depreciation expense $20000
Long term Asset account $20000
Adjusted Depreciation expense account $20000
Accumulated Depreciation expense account $20000


3) Depreciation cost for 4 years = 4*28600= 114,4003) If the Bailand has ignored the residual value, the depreciation cos = 286,000/10 = 28,600

Difference between the two methods = 114400-112000 = $2,400

Which means $2400 need to be reduced from the remaining amount to be paid for the depreciation.

ie., remaing amount for depreciation = (280000-112000-2400)/6 = 165600/6 = $27,600

Journal Entry Credit Debit
Accumulated Depreciation expense Account $28,600
Long term Asset account $28,600
Accumulated Depreciation expense account $1000
Adjusted Depreciation expense $1000
Adjusted Depreciation expense account $1000
Long term Asset account $1000